# Towards Efficient Transferable Preemptive Adversarial Defense

### v3

@@ Abstract
Deep learning technology has brought convenience and advanced developments but has become untrustworthy because of its sensitivity to inconspicuous perturbations (i.e., adversarial attacks). Attackers may utilize this sensitivity to manipulate predictions. To defend against such attacks, we have devised a proactive strategy for “attacking” the medias before it is attacked by the third party, so that when the protected medias are further attacked, the adversarial perturbations are automatically neutralized. This strategy, dubbed Fast Preemption, provides an efficient transferable preemptive defense by using different models for labeling inputs and learning crucial features. A forward-backward cascade learning algorithm is used to compute protective perturbations, starting with forward propagation optimization to achieve rapid convergence, followed by iterative backward propagation learning to alleviate overfitting. This strategy offers state-of-the-art transferability and protection across various systems. With the running of only three steps, our Fast Preemption framework outperforms benchmark training-time, test-time, and preemptive adversarial defenses. We have also devised the first to our knowledge effective white-box adaptive reversion attack and demonstrate that the protection added by our defense strategy is irreversible unless the backbone model, algorithm, and settings are fully compromised. This work provides a new direction to developing proactive defenses against adversarial attacks. The proposed methodology will be made available on GitHub.
Index Terms:
Transferable preemptive adversarial defense, adaptive reversion attack, proactive defense, adversarial attack.
I Introduction
Deep learning technologies have been demonstrated to be sensitive to inconspicuous perturbations (i.e., adversarial attacks) [1]. An adversarial attacker can utilize this weakness to manipulate medias such as images, audios, and videos. For example, the attacker can change the identity of a face image to that of the attacker without changing the looking, enabling the attacker to illegally access (i.e., hack into) the face authentication system [2, 3]. The attacker can also add slight noise on traffic signs, possibly causing an accident during autonomous driving [4]. Defending against such attacks is a critical but unsolved challenge. Passive defenses, on the one hand, generally sacrifice classification accuracy for resistance against threats [5, 6, 7, 8, 9]. Proactive defenses, on the other hand, do not achieve satisfactory time efficiency when running protection in real-time applications. Additionally, these proactive defenses typically require paired modules for protection and inference, so their protection is not transferable to other systems [10, 11]. Furthermore, if a provided model is used, the user may concern if backdoor exists in this model [12, 13].
@@ I Introduction
To address these problems, we devise a proactive strategy, called Fast Preemption, as illustrated in Figure 1. Fast Preemption preemptively “attacks” (manipulates for enhancement) the medias from the user side before being attacked by the third party. Specifically, Fast Preemption learns robust features and predicts possible adversarial perturbations, then enhances these robust features and reverses the predicted perturbations for purpose of protection. In this case, when these protected medias are further attacked, these adversarial perturbations will be automatically neutralized. This strategy has three distinct advantages. (i) A proactive preemptive defense provides better resistance against threats than a passive one without degrading accuracy, whereas Fast Preemption requires no additional operation for the classification system. (ii) Fast Preemption allows plug-and-play implementation because it does not require paired modules, alleviating the security and privacy concerns through a backdoor when using another provider’s system. (iii) From the attacker’s perspective, once a training- or test-time defense (typically integrated at the classification model) is breached, all inputs are affected. In contrast, with this preemptive defense, only a single user is affected as users are likely to be using different settings. Therefore, Fast Preemption meets four basic requirements:
@@
As illustrated in Figure 2, the Fast Preemption framework includes three components: a classifier model for labeling inputs, a backbone model for learning features, and the forward-backward (Fore-Back) cascade learning algorithm for computing protective perturbations. The classifier determines the correctness of the optimization direction. The backbone model greatly determines transferability and time cost. However, the best backbone model must be adversarially trained [5, 14], and its clean accuracy is typically sacrificed for robustness against threats. The Fast Preemption framework uses different models for the classifier and backbone, thereby optimizing both tasks. Additionally, as we demonstrate, the state-of-the-art (SOTA) preemptive defense [10] cannot achieve satisfactory time cost for efficient user interaction. We thus developed the Fore-Back cascade learning algorithm combining forward propagation optimization to achieve rapid convergence and backward propagation learning to alleviate overfitting.
We conducted evaluation experiments and ablation studies on two datasets [15, 16] across four models [5, 6, 7, 8], defending against two typical adversarial attacks [17, 18] in three settings. The results demonstrate the superior performance of the Fast Preemption strategy in terms of clean and robust accuracies, transferability, and time cost compared with those of benchmark training-time, test-time, and preemptive adversarial defenses [7, 8, 9, 10, 11]. For example, in an end-to-end comparison with the SOTA preemptive defense [10] on the ImageNet dataset [16], Fast Preemption attained 1.5% and 5.6% increases in terms of the clean and robust accuracies, respectively, against AutoAttack (Standard) [18] and was 33 times faster.
Shannon insightfully claimed that “one ought to design systems under the assumption that the enemy will immediately gain full familiarity with them [19].” On the basis of that assumption, we designed the first effective white-box adaptive reversion attack, dubbed Preemptive Reversion, for evaluating the reliability of preemptive defenses, i.e., whether they are reversible. As illustrated in Figure 3, Preemptive Reversion conducts the secondary preemptive defense upon the protected example and then reverses these new perturbations because, in the same setting, the secondary defense offers mostly the same perturbations as the first defense. Along with adaptive reversion, we propose using a more comprehensive evaluation protocol, one that requires a decrease in the clean accuracy after the instance is preemptively defended (which is ensured by selecting a poor classifier). If clean accuracy increases after reversion, the reversion succeeded. If it decreases after reversion (due to distortion), reversion failed.
@@
  • We present a new preemptive defense strategy (Fast Preemption) providing more transferable protection without requiring ground truths or requiring users to use a paired module. It is the first to use different models for the classifier and backbone.
  • We present a forward-backward (Fore-Back) cascade learning algorithm that speeds up preemptive defense for efficient user interaction.
  • We present the first effective adaptive reversion attack (Preemptive Reversion) for evaluating the reliability of preemptive defenses in a white-box setting and a more comprehensive protocol to evaluate whether reversion succeeds.
Additionally, we will release a new toolbox upon acceptance that enables plug-and-play implementation with multiple options for each component. This toolbox can also integrate benchmark preemptive defenses.
II Related Work
Training-time defenses produce robust deep learning models that classify adversarial examples (AEs) correctly. The earliest mechanisms, e.g., adversarial training [1], incorporated various types of AEs along with their ground-truth labels in the training dataset. To improve the efficiency of training on large-scale datasets, Shafahi et al. [20] devised accelerated adversarial training in which backward-pass computations are reused. This entails curating a large number of specific real-world samples; Sehwag et al. [21] circumvented this problem by using additional data from proxy distributions learned by advanced generative models for adversarial training. Gowal et al. [5] demonstrated that generative models trained solely on the original dataset can be leveraged to artificially increase the training dataset’s size and improve adversarial robustness against $L_{p}$-norm perturbations. Debenedetti et al. [6] proposed using a custom recipe of vision transformers for adversarial training in which weaker data augmentation is used with warm-up in the augmentation intensity. Inspired by denoising diffusion models [22], Wang et al. [23] conducted adversarial training using data generated by an elucidating diffusion model [24]. Dong et al. [25] introduced robust examples (REs) generated by a preemptive defense into the adversarial training and demonstrated alleviation of clean accuracy deterioration. The current SOTA robust model, devised by Peng et al. [7], follows a suite of generalizable robust architectural design principles: (i) an optimal range for depth and width configurations; (ii) a preference for a convolutional rather than patchify stem stage, and (iii) a robust residual block design featuring squeeze and excitation blocks and nonparametric, smooth activation functions.
@@ II Related Work
Training-time defenses have three shortcomings. First, they require vast amounts of data for training, and the data quality and diversity are crucial to robustness. Second, robust models can be counterattacked. Third, training-time defenses typically reduce clean accuracy in favor of robust accuracy, even the one proposed by Dong et al. [25], which was designed to circumvent this problem.
Test-time defenses combine detection and purification. Detection models are typically trained on AEs to enable the model to distinguish them from benign inputs. Among recent studies, Cohen et al. [26] devised a k-nearest neighbor model for ranking the most supportive training samples for any given validation example in training a detector. Zuo and Zeng [27] proposed using a binary classifier, erase-and-restore, that uses the intriguing sensitivity of $L_{2}$-norm attacks by randomly erasing certain pixels in an $L_{2}$-norm AE and then restoring them with an inpainting technique. Liang et al. [28] detected AEs by comparing the classification results of a given sample and a denoised version that was generated through adaptive noise reduction combining scalar quantization and a smoothing spatial filter. Most recently, Hickling et al. [29] devised a detector using explainable deep reinforcement learning for uncrewed aerial vehicle agents. However, detection of unseen attacks is challenging for all existing detectors because they are trained on AEs from known attacks. Additionally, detection always introduces errors due to false classification of benign images as AEs, thus reducing clean accuracy.
In contrast, adversarial purification aims to purify AEs by eliminating or invalidating perturbations. Earlier studies used image distortion for purification. For instance, Wang et al. [30] distorted images by randomly nullifying pixels. Guo et al. [31] preprocessed inputs with a nondifferentiable transformation. Raff et al. [32] combined a large number of individually weak defenses into a single randomized transformation to build a strong defense. Lately, more practical approaches generate benign replacements for inputs regardless of whether they are malicious. Ren et al. [33] trained a generative model using autoencoder technology [34]. They claimed the existence of an immune space, in which perturbations have fewer adverse effects, and they inactivated adversarial perturbations by restricting them to that subspace. Nie et al. [9] devised the DiffPure benchmark, which uses denoising diffusion models [22]. DiffPure replaces perturbations with Gaussian noise and then restores the clean image by reversing the diffusion process. The major limitation of purification is that the difference between the original and generated examples degrades clean accuracy.
@@
III Preemptive Defense Strategy: Fast Preemption
@@
III-A Problem Definition
Adversarial attacks and preemptive defenses both manipulate the input by adding subtle perturbations. Specifically, let $x$ and $y$ denote the original image and its ground-truth label, respectively, and let $f_{v}$ denote the victim model. Let $x^{\prime}$ denote the manipulated example (AE or RE), formulated as
@@ III-A Problem Definition
where $||\cdot||_{p}$ represents the $L_{2}$- or $L_{\infty}$-norm and $\epsilon$ is the setting of the maximum distance. Therefore, adversarial attacks and preemptive defenses aim to solve different optimization problems (as explained below) by changing $\delta$ with the inputs of $x$, $y$, and $f_{v}$.
III-A1 Adversarial Attack
@@
III-A2 Preemptive Defense
A preemptive defense aims to generate an RE that will make Equation (2) less successful. Specifically, with the preemptive defense strategy, “backward propagation optimization” is used to reverse the adversarial attack perturbations [10]. In other words, the optimization of Equation (2) remains, but output $\delta$ is reversed. The backward propagation optimization is formulated as
@@ III-A2 Preemptive Defense
In contrast to backward propagation optimization, we newly define “forward propagation optimization” for the preemptive defense strategy. It directly reverses the optimization direction of Equation (2) from maximization to minimization, which in turn enhances the model’s confidence in accurately classifying the input according to its true label:
@@
Both the forward and backward optimization learning algorithms will be further addressed in Section III-C.
III-B Framework
As illustrated in Figure 2, the Fast Preemption framework includes three components.
Classifier: The correctness of the input labels determines the correct optimization direction for defense. However, for efficient user interaction, it is impractical to label each input manually. Classifier $f_{c}$ labels these inputs automatically. Therefore, $y$ in Equations (3) and (4) can be replaced by $f_{c}(x)$. The framework is integrated with user application before being attacked, so $f_{c}$ does not have to show any transferability or resistance against threats, but it must offer the best classification accuracy. Processing images on the basis of the ground truth can be regarded as using a perfect classifier.
Backbone Model: With Fast Preemption, the input image is iteratively optimized on the basis of feedback from backbone model $f_{b}$, which may differ from victim model $f_{v}$ in Equations (3) and (4). Therefore, due to its complexity, $f_{b}$ determines the transferable clean and robust accuracies and time cost. To achieve better performance, the backbone model $f_{b}$ must be adversarially trained so it can learn more crucial features instead of random pixels. The limitations of adversarially trained models (i.e., training-time defenses), such as the degradation of clean accuracy, are alleviated by the well-performing nonrobust classifier $f_{c}$ in our framework. As a result, the preemptive defense optimizes the inputs towards the correct labels. Therefore, two different models, classifier $f_{c}$ and backbone $f_{b}$, are integrated into the framework.
Fore-Back Cascade Learning Algorithm: To meet our framework, the optimization problems formulated as Equations (3) and (4) can be reformulated as
@@ III-B Framework
III-C Forward-Backward (Fore-Back) Cascade Learning
Algorithm
The Fore-Back cascade learning algorithm is used to compute protective perturbations. The input is first moved away from the classification boundaries of all incorrect classes so that the manipulated input will be easier to be classified as the ground truth. We define this as forward propagation ($\rightarrow$). Forward propagation can quickly attain convergence (less time cost) but causes overfitting on the backbone model so that the defense is less transferable to other victim models. Backward propagation optimization ($\leftarrow$) is thus conducted following forward propagation. Backward propagation reverses the adversarial perturbations to make them protective perturbations, which effectively alleviates overfitting and thereby improves transferability. In summary, the forward and backward propagation both enhance robustness, but in different optimization directions.
III-C1 Loss Function
From Equations (5) and (6), the loss function of our Fore-Back cascade learning algorithm is the CE loss:
@@ III-C1 Loss Function
III-C2 Preliminary - Backward Propagation
The use of backward propagation for preemptive defense was introduced in the SOTA preemptive defense strategy [10]. At each step of that strategy, a PGD attack [17] is run first as $x_{1}\rightarrow x_{1}^{a}$ (Figure 2), and then the adversarial perturbations are reversed in accordance with the learning rate and thereby forms protective perturbations, as $x_{1}\rightarrow x_{2}$.

### v4

# Fast Preemption: Forward-Backward Cascade Learning for Efficient and Transferable Proactive Adversarial Defense

@@ Abstract
Deep learning technology has brought convenience and advanced developments but has become untrustworthy due to its sensitivity to adversarial attacks. Attackers may utilize this sensitivity to manipulate predictions. To defend against such attacks, existing anti-adversarial methods typically counteract adversarial perturbations post-attack, while we have devised a proactive strategy that preempts by safeguarding media upfront, effectively neutralizing potential adversarial effects before the third-party attacks occur. This strategy, dubbed Fast Preemption, provides an efficient transferable preemptive defense by using different models for labeling inputs and learning crucial features. A forward-backward cascade learning algorithm is used to compute protective perturbations, starting with forward propagation optimization to achieve rapid convergence, followed by iterative backward propagation learning to alleviate overfitting. This strategy offers state-of-the-art transferability and protection across various systems. With the running of only three steps, our Fast Preemption framework outperforms benchmark training-time, test-time, and preemptive adversarial defenses. We have also devised the first, to our knowledge, effective white-box adaptive reversion attack and demonstrate that the protection added by our defense strategy is irreversible unless the backbone model, algorithm, and settings are fully compromised. This work provides a new direction to developing proactive defenses against adversarial attacks.
1 Introduction
Deep learning technologies have been demonstrated to be sensitive to adversarial attacks [16]. An adversarial attacker can utilize this weakness to manipulate medias such as images, audios, and videos. For example, the attacker can change the identity of a face image to that of the attacker without changing the looking, enabling the attacker to illegally access the face authentication system [43]. The attacker can also add slight noise on traffic signs, possibly causing an accident during autonomous driving [13]. Defending against such attacks is a critical but unsolved challenge. Passive defenses [5, 26, 31, 20], on the one hand, trade off classification accuracy for resistance against threats [46], but may be vulnerable to counterattacks [6, 42]. Proactive defenses [29, 15], on the other hand, do not achieve satisfactory time efficiency when running protection in real-time applications. Additionally, these proactive defenses typically require paired modules for protection and inference, so their protection is untransferable to other systems. Furthermore, if a provided model is used, the user may concern if backdoor exists in this model [35].
@@ 1 Introduction
To address these problems, we devise a proactive strategy, called Fast Preemption, as illustrated in Fig. 1. Fast Preemption preemptively “attacks” (manipulates for enhancement) the medias from the user side before being attacked by the third party. Specifically, Fast Preemption learns robust features and predicts possible adversarial perturbations, then enhances these robust features and reverses the predicted perturbations for purpose of protection. In this case, when these protected medias are further attacked, these adversarial perturbations will be automatically neutralized. This strategy has three distinct advantages. (i) A proactive preemptive defense provides better resistance against threats than a passive one without degrading accuracy, whereas Fast Preemption requires no additional operation for the classification system. (ii) Fast Preemption allows plug-and-play implementation because it does not require paired modules, alleviating the security and privacy concerns through a backdoor when using another provider’s system [14, 35]. (iii) From the attacker’s perspective, once a training- or test-time defense (typically integrated within a classification model) is breached, all inputs are affected. In contrast, with this preemptive defense, only a single user is affected as users are likely to be using different settings. Therefore, Fast Preemption meets four basic requirements:
@@
We propose the Fast Preemption defense includes three components: a classifier model for automatically labeling inputs, a backbone model for learning features (differ from the classifier), and the forward-backward (Fore-Back) cascade learning algorithm for efficiently computing transferable protective perturbations. We conducted evaluation on two datasets [24, 1] across six models [17, 5, 32, 26, 29, 11], defending against two typical adversarial attacks [28, 9]. The results demonstrate the superior performance of the Fast Preemption strategy in terms of clean and robust accuracies, transferability, and time cost compared with those of benchmark training-time, test-time, and preemptive adversarial defenses [5, 26, 31, 29, 15]. For example, in an end-to-end comparison with the state-of-the-art (SOTA) preemptive defense [29] on the ImageNet dataset [1], Fast Preemption attained 1.5% and 5.6% increases in terms of the clean and robust accuracies, respectively, against AutoAttack (Standard) [9] and was 33 times faster.
Shannon insightfully claimed that “One ought to design systems under the assumption that the enemy will immediately gain full familiarity with them [39].” On the basis of that assumption, we designed the first effective white-box adaptive reversion attack, dubbed Preemptive Reversion, for evaluating the reliability of preemptive defenses, i.e., whether they are reversible. Preemptive Reversion conducts the secondary preemptive defense upon the protected example and then reverses these new perturbations because, in the same setting, the secondary defense offers mostly the same perturbations as the first defense. Along with adaptive reversion, we propose using a more comprehensive evaluation protocol, one that requires a decrease in the clean accuracy after the instance is preemptively defended (which is ensured by selecting a poor classifier). If clean accuracy increases after reversion, the reversion succeeded. If it decreases after reversion (due to distortion), reversion failed.
@@
We summarize our contributions as follows: (i) We present a new preemptive defense strategy (Fast Preemption) providing more transferable protection without requiring ground truths or requiring users to use a paired module. It is the first to use different models for the classifier and backbone. (ii) We present a forward-backward (Fore-Back) cascade learning algorithm that speeds up preemptive defense for efficient user interaction. (iii) We present the first effective adaptive reversion attack (Preemptive Reversion) for evaluating the reliability of preemptive defenses in a white-box setting and a more comprehensive protocol to evaluate whether reversion succeeds.
2 Related Work
Existing adversarial defenses can be categorized into three primary types: training-time, test-time, and preemptive defenses. Appendix A provides a more detailed review.
@@ 2 Related Work
Training-time defenses produce robust deep learning models that may correctly classify adversarial examples (AEs) as their ground truth labels [16, 38, 37, 17, 11, 45, 12, 32, 46, 5, 26, 40]. However, they have two shortcomings. First, they require vast amounts of data for training, and the data quality and diversity are crucial to robustness. Second, training-time defenses typically reduce clean accuracy in favor of robust accuracy.
Test-time defenses are divided to detection and purification. Detection models are typically trained on AEs to enable the model to distinguish them from benign inputs [7, 47, 25, 20]. However, detection of unseen attacks is challenging for all existing detectors because they are trained on AEs from known attacks. Additionally, detection always introduces errors due to false classification of benign images as AEs, thus reducing clean accuracy. In contrast, adversarial purification aims to purify AEs by eliminating or invalidating perturbations. Earlier studies used image distortion for purification [44, 18, 33]. Lately, more practical approaches generate benign replacements for inputs regardless of whether they are malicious [34, 31]. The major limitation of purification is that the difference between the original and generated examples degrades clean accuracy, and the possibility of being counterattacked [6, 42].
Preemptive defenses seek to generate robust examples (REs) that replace the original images, which are subsequently discarded [36, 29, 15]. These REs are inherently more difficult to manipulate, thereby providing resistance against adversarial attacks. However, such approaches face significant limitations, including high computational costs, limited transferability across models, and, in some cases, lower robustness compared to other defense mechanisms.
Lastly, Alfarra et al. [2] proposed a test-time defense that integrates the preemptive defense technique as an anti-adversary layer. However, the use of preemptive defense techniques for test-time defenses is less practical because test-time defenses do not know whether inputs are malicious, so an anti-adversary layer may fail to resist certain attacks because it relies on a label that was predicted from the original classifier and may initially be incorrect.
@@
3 Preemptive Defense: Fast Preemption
@@
3.1 Problem Definition
Adversarial attacks and preemptive defenses both manipulate inputs by adding subtle perturbations. Specifically, let $x$ and $y$ denote the original image and its ground-truth label, respectively, and let $f_{v}$ denote the victim model. Let $x^{\prime}$ denote the manipulated example (AE or RE), formulated as
@@ 3.1 Problem Definition
where $\delta$ denotes the perturbations, $||\cdot||_{p}$ represents the $L_{2}$- or $L_{\infty}$-norm and $\epsilon$ is the setting of the maximum distance. Therefore, adversarial attacks and preemptive defenses aim to solve different optimization problems (as explained below) by changing $\delta$ with the inputs of $x$, $y$, and $f_{v}$.
3.1.1 Adversarial Attack
@@
3.1.2 Preemptive Defense
A preemptive defense aims to generate REs those will make Eq. 2 less successful. With the preemptive defense strategy, “backward propagation optimization” is used to reverse the adversarial attack perturbations [29]. In other words, the optimization of Eq. 2 remains, but output $\delta$ is reversed. The backward propagation optimization is formulated as
@@ 3.1.2 Preemptive Defense
In contrast to backward propagation optimization, we newly define “forward propagation optimization” for the preemptive defense strategy. It directly reverses the optimization direction of Eq. 2 from maximization to minimization, which in turn enhances the model’s confidence in accurately classifying the input according to its true label:
@@
3.2 Framework
As illustrated in Fig. 2, the Fast Preemption framework includes three components.
Classifier: Correct input labels determines the correct optimization direction for defense. However, for efficient user interaction, it is impractical to label each input manually. Classifier $f_{c}$ labels these inputs automatically. Therefore, $y$ in Eq. 3 and Eq. 4 can be replaced by $f_{c}(x)$. The framework is integrated with user application before being attacked, so $f_{c}$ does not have to show any transferability or resistance against threats, but it must offer the best classification accuracy. Processing images with ground truths can be regarded as using a perfect classifier.
Backbone Model: With Fast Preemption, the input image is iteratively optimized on backbone model $f_{b}$, which may differ from victim model $f_{v}$ in Eq. 3 and Eq. 4. Therefore, due to its complexity, $f_{b}$ determines the transferable clean and robust accuracies and time cost. To achieve better performance, the backbone model $f_{b}$ must be adversarially trained so it can learn more crucial features instead of random pixels. The limitations of adversarially trained models (i.e., training-time defenses), such as the degradation of clean accuracy, are alleviated by the well-performing nonrobust classifier $f_{c}$ in our framework. As a result, the preemptive defense optimizes the inputs towards the correct labels. Therefore, two different models, classifier $f_{c}$ and backbone $f_{b}$, are integrated into the framework.
Fore-Back Cascade Learning Algorithm: To meet our framework, the optimization problems formulated as Eq. 3 and Eq. 4 can be reformulated as
@@ 3.2 Framework
3.3 Forward-Backward Cascade Learning
The Fore-Back cascade learning algorithm is used to compute protective perturbations. The input is first moved away from the classification boundaries of all incorrect classes so that the manipulated input will be easier to be classified as the ground truth. We define this as forward propagation ($\rightarrow$). Forward propagation can quickly attain convergence (less time cost) but causes overfitting on the backbone model so that the defense is less transferable to other victim models. Backward propagation optimization ($\leftarrow$) is thus conducted following forward propagation. Backward propagation reverses the adversarial perturbations to make them protective perturbations, which effectively alleviates overfitting and thereby improves transferability. In summary, the forward and backward propagation both enhance robustness, but in different optimization directions. The complete algorithm is presented in Algorithm 1 of Appendix B.
3.3.1 Loss Function
From Eq. 5 and Eq. 6, the loss function of our Fore-Back cascade learning algorithm is the CE loss:
@@ 3.3.1 Loss Function
3.3.2 Preliminary - Backward Propagation
The use of backward propagation for preemptive defense was introduced in the SOTA preemptive defense strategy [29]. At each step of that strategy, a projected gradient descent (PGD) attack [28] is run first as $x_{1}\rightarrow x_{1}^{a}$ (Fig. 2), and then the adversarial perturbations are reversed in accordance with the learning rate and thereby forms protective perturbations, as $x_{1}\rightarrow x_{2}$.